2. If p, q, r and s are in A.P. then r – q is
(a) s – p
(b) s – q
(c) s – r
(d) none of these

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Answer: c Explaination:Reason: Since p, q, r, s are in A.P. ∴ (q – p) = (r – q) = (s – r) = d (common difference)